# Learn while Unlearn: An Iterative Unlearning Framework for Generative Language Models

### v1

@@ 3.6 Training
${}\mathcal{L}=\mathcal{L}_{fgt}+\alpha\mathcal{L}_{lrn}+\beta\mathcal{L}_{kl}$ (5)
where $\alpha,\beta>0$ are hyper-parameters that modulate the salience of different objectives.
@@ 4 Experiments
In this section, we first describe the datasets utilized for training and evaluation, and introduce the metrics employed to assess performance. Next, we present the baseline methods used for comparison with our proposed method. We then detail the configuration of our proposed method. Finally, we will present the experimental results and analysis in detail.
@@ 4.1 Datasets
To evaluate the method’s learning and unlearning capabilities, we selected two types of datasets: Target Datasets, which assess unlearning ability, and Downstream Dataset, which evaluate the original capabilities of the models.
Target Dataset
The Pile corpus (825GB) is a large dataset constructed from 22 diverse high-quality subsets, many of which derive from academic or professional sources (e.g. books, open source code) [20]. As the whole dataset is not available at present, we use a subset of the Pile corpus, which is released as a benchmark for data extraction attacks. [^2] https://github.com/google-research/lm-extraction-benchmark Designed to be easy-to-extract, the subset contains 15,000 samples, randomly sampled from the Pile training dataset. Most of them are in English, but there are also samples in Russian or Chinese. Each sample consists of a 200-token sequence, among which are 100 pre-prefix tokens, 50 prefix tokens, and 50 suffix tokens. We only use the prefix and suffix tokens in our experiments.
Downstream Dataset
To assess the general performance of the LMs subsequent to the process of unlearning, a diverse array of downstream tasks is employed. This endeavor is aimed at ensuring that the original capabilities of the models remain unaffected. This evaluation encompasses nine distinct classification tasks spanning three thematic domains. Specifically, these domains include linguistic reasoning tasks such as Hellaswag [58] and Lambada [39], as well as assessments of commonsense reasoning through Winogrande [46] and COPA [22]. Additionally, scientific reasoning abilities are evaluated through tasks such as ARC-Easy [15], ARC-Challenge [15], Piqa [5], MathQA [2], and PubmedQA [28]. Furthermore, four dialogue tasks, namely Wizard of Wikipedia [17], Empathetic Dialogues [45], Blended Skill Talk [47], and Wizard of Internet [32], are used to gauge the model’s proficiency in generating coherent responses. In addition, we measure the perplexity of the unlearned models on the validation set of Pile and Wikitext. As per Jang et al. [27], we use the test set for Lambada and the validation set for the remaining tasks.
@@ 4.2 Metrics
As stated in Section 4.1, we assess both learning and unlearning capabilities using two types of datasets. For unlearning ability, we follow Jang et al. [27] and examine the forgetting effect on the target unlearning data by EL and MA. EL and MA are used as the standard of terminating the training process, which makes them not suitable to be the evaluation metrics alone. Meanwhile, as mentioned in Section 3.5, we employ the BERTScore and BLEU to measure forgetting during IUR. For evaluating learning ability, we first utilize the metrics provided in the downstream datasets to measure performance. We then test the information entropy to ensure the result is coherent. Furthermore, we use GPT-4 to evaluate the text generated by the unlearned models. The following are the details of the metrics used.
Extraction Likelihood (EL) is introduced by Jang et al. [27] to measure the average success rate of varying extraction attacks quantified via getting the n-gram overlap of generated and target token sequences. It is computed by the following equation:
$\mathrm{EL}_{n}(\boldsymbol{x})=\frac{\sum_{t=1}^{T-n}\mathrm{OVERLAP}_{n}(f_{\theta}(x_{<t}),x_{\geq t})}{T-n}$ (6)
$\mathrm{OVERLAP}_{n}(\boldsymbol{a},\boldsymbol{b})=\frac{\sum_{c\in ng(\boldsymbol{a})}\mathbbm{1}\{c\in ng(\boldsymbol{b})\}}{|ng(\boldsymbol{a})|}$ (7)
Memorization Accuracy (MA) [49] quantifies how much model $f_{\theta}$ has memorized the given token sequences, which is defined as follows:
$\mathrm{MA}(\boldsymbol{x})=\frac{\sum_{t=1}^{T-1}\mathbbm{1}\{\mathrm{argmax}(P_{\theta}(\cdot|x_{<t})=x_{t}\}}{T-1}$ (8)
@@
$H(x)=-\sum_{i=1}^{n}p_{i}\log_{2}p_{i}$ (9)
GPT Evaluation uses GPT-4 [1] to evaluate the unlearned models in two perspectives: whether the model generates text without prior knowledge of key information in the referenced target data and whether the generated sequences are coherent. The prompts can be found in Appendix C.
@@ 4.3 Baseline Methods
Our experiments use the GPT-Neo model family (125M, 1.3B, 2.7B) [6], which is pre-trained on the Pile corpus. We use the Opt model family (125M, 1.3B, 2.7B) [59], which is pre-trained on a deduplicated version of the Pile as well as other corpus. Following Jang et al. [27], Opt serve as our baseline method for deduplication. Since the deduplicated version of GPT-Neo by Kandpal et al. [29] is not publicly available, we include KUMPR [27] as a second baseline method on GPT-Neo models to show the effectiveness of our proposed method. We follow the same training and evaluation procedure as Jang et al. [27] for the baseline methods.
@@ 4.4 Configurations
For each model size, we execute five runs of the methods, each targeting a dataset of 128 samples. Utilizing the all-MiniLM-L6-v2 model, we employ KNN to identify samples within the Pile subset, selecting the nearest one as the learning target. To determine the optimal hyperparameters for our proposed approach, we conduct multiple experiments involving varying learning rates as well as different weights assigned to $\mathcal{L}_{fgt}$, $\mathcal{L}_{lrn}$, and $\mathcal{L}_{nor}$. The model is optimized by Adam [31] with a learning rate of $65e-$ 6, and $\alpha=0.5,\beta=1.0$. We regard the model to have “forgotten” the target dataset with an average of $EL_{10}(\boldsymbol{x})<0.0499$ and $MA(\boldsymbol{x})<0.5994$. The filtering thresholds during iteration are $a=0.3$ and $b=0.01$ as introduced in Section 3.5.
@@
Table 1: Main Results showing the average of 5 random samples. Cls Avg. denotes the average accuracy of the 9 classification datasets, and Dia Avg. denotes the average F1 score of the 4 dialogue datasets. The best comparable performances are bolded and second best underlined.
@@ 4.5 Main Results
The results of all methods are summarized in Table 1. Overall, our method consistently achieves the best or second-best performance across all metrics when compared to the baselines. Specifically, regarding unlearning ability (measured by EL, MA, and BERT), our method is second only to KUMPR, which uses these metrics as halting signals, and significantly outperforms the original model without unlearning methods. In terms of learning ability, our method attains near-optimal performance on most metrics and surpasses KUMPR across all metrics. Additionally, we discover some interesting phenomena as follows:
First, the impact of unlearning becomes more pronounced with larger models, indicating a higher propensity for larger models to memorize sensitive information, which our method effectively mitigates. Second, while the KUMPR approach significantly forgets sensitive information, it also impairs the model’s normal ability (e.g, PPL). In contrast, our ICU approach preserves the model’s essential linguistic capabilities while erasing sensitive information, underscoring the necessity of our methods. The full results of each sample can be found in Appendix A.
@@ 4.6 Ablation Study
As discussed in Section 3.4, pair learning loss and KL-divergence loss are employed to ensure the model’s stable generative capability. To assess the impact of these losses, Figure 3 presents the performance after removing them respectively. The results indicate that both losses enhance learning performance, affirming their role in preserving the model’s original generative capacity. Furthermore, the KL-divergence loss has a more significant effect, suggesting that aligning the model’s output distribution with the original model is crucial for maintaining its generative capacity.
@@ 4.7 Parameter Sensitivity
5 Case Study
@@
To provide a clearer comparison of our methods, we present a case study demonstrating the balance between learning and unlearning. As illustrated in Figure 4, the Reference includes sensitive information such as “@ThingsExpo 2016”. Before unlearning, the original models (e.g., GPT-Neo and Opt) retain this information and can reproduce it using the corresponding prefix. When the KUMPR method is applied, the models lose their original conversational abilities and repetitively output the word “the”. Our approach, in contrast, effectively forgets the sensitive information while simultaneously learning the correct outputs from paired data. This ensures the model retains its generative capabilities, highlighting the significance of our Iterative Contrastive Unlearning framework. Other examples can be found in Appendix D.
Figure 4: An example for case study.
6 Conclusion
@@
In this paper, we explored the motivated direction of machine unlearning for Generative Language Models. We first analyzed the challenges in this topic, and further proposed an Iterative Contrastive Unlearning framework (ICU). Specifically, we extend the Knowledge Unlearning Induction with Contrastive Learning Enhancement, which trains the model with selected similar data. In addition, we introduce Iterative Unlearning Refinement to prevent further unlearning on discarded information, preserving the model’s ability adaptively. Finally, extensive experiments of models in three different sizes demonstrate the effectiveness of our proposed method. We hope our work will lead to more future studies.
7 Limitation
Generative language models typically possess a considerable number of parameters, necessitating re-tuning during machine unlearning. As the volume of data to be forgotten expands, the number of iterations required for unlearning also increases, leading to a notable escalation in computational expenses. Moreover, the process of unlearning may necessitate the management of sensitive user information or personal data, potentially giving rise to privacy and security considerations. It is imperative to prioritize safeguarding user privacy and adhering to applicable laws and regulations throughout the unlearning procedure.
Appendix A Full Results
The full results of the individual runs are shown in Table 3 and 4. Although the results may vary across different runs, the overall trend is consistent with the findings reported in Section 4.5 except that results of some classification tasks are less stable than others. Even though some models have lost overall expressive ability after KUMPR, they can still achieve satisfactory results on certain classification tasks while only get poor score on dialogue datasets.
Table 3: Full Results of 5 individual runs.
Neo 125M 51.9 76.8 70.3 4.139 37.6 28.2 51.5 62.0 46.0 22.4 63.4 22.5 57.6 43.5
+ KUMPR 125M 0.1 22.0 26.5 0.863 2.0 26.9 51.1 57.0 36.0 22.1 56.9 21.3 41.0 34.9
0.3 24.0 33.3 0.707 44.8 27.3 52.6 49.0 38.9 21.1 59.6 22.5 57.6 41.5
1.2 9.4 28.9 0.546 0.0 26.0 50.6 56.0 30.7 20.1 52.0 20.7 32.4 32.1
0.4 18.2 30.8 0.327 0.1 26.6 49.9 56.0 32.8 21.4 55.5 21.2 32.4 32.9
1.5 22.0 29.1 1.119 0.5 26.8 50.7 57.0 33.7 21.1 26.7 21.1 40.6 34.2
+ ours 125M 3.6 53.3 52.8 3.848 37.3 28.3 51.1 62.0 45.3 23.1 63.2 22.6 57.6 43.4
4.8 52.3 53.0 3.670 44.4 28.4 53.1 60.0 43.9 20.7 62.8 22.6 57.6 43.7
4.4 56.3 51.8 3.770 34.2 28.3 52.2 61.0 43.9 22.4 63.3 21.7 57.8 42.8
3.9 59.0 54.8 3.740 36.1 28.3 51.9 62.0 43.7 20.7 62.7 22.5 57.6 42.8
5.0 56.9 54.3 4.137 43.1 28.3 52.7 61.0 44.9 22.1 62.7 22.2 57.6 43.8
Neo 1.3B 98.2 92.3 86.3 4.640 57.4 37.0 54.9 70.0 56.5 25.8 70.3 22.0 53.8 49.7
+ KUMPR 1.3B 0.0 0.0 37.8 2.958 0.0 25.7 50.8 51.0 25.4 17.4 53.2 18.2 57.6 33.3
0.6 7.0 21.7 0.216 0.0 25.7 50.4 55.0 28.2 19.4 53.3 21.3 57.6 34.5
0.7 10.5 28.1 0.471 0.0 25.9 48.7 52.0 27.9 19.4 52.7 20.6 57.6 33.9
1.0 9.8 22.5 0.324 0.0 25.9 50.3 55.0 27.7 19.4 52.4 19.7 57.6 34.2
1.6 13.2 22.7 0.115 0.0 25.8 49.8 50.0 30.0 22.4 52.3 20.4 57.6 34.3
+ ours 1.3B 4.5 53.3 53.5 4.218 54.3 37.1 54.8 70.0 55.6 25.4 69.8 21.3 49.4 48.6
4.8 53.8 57.7 4.069 54.3 37.2 55.3 69.0 56.3 25.8 70.2 21.8 50.2 48.9
4.5 53.7 53.8 4.488 55.7 37.6 55.2 68.0 56.0 26.4 70.4 21.4 50.0 49.0
4.8 41.0 44.2 2.440 54.6 37.3 55.1 70.0 56.8 25.8 69.9 21.5 49.8 49.0
4.7 54.5 54.4 4.286 57.5 37.3 54.9 67.0 56.7 25.8 70.1 21.5 50.5 49.3
Neo 2.7B 96.7 93.7 90.2 4.719 62.3 40.8 56.6 75.0 59.5 26.1 73.0 21.3 57.0 52.4
+ KUMPR 2.7B 1.1 29.4 28.2 0.603 7.9 26.3 48.2 53.0 30.3 19.1 53.4 20.2 57.2 35.1
3.0 25.4 25.1 0.304 0.8 25.9 50.4 57.0 30.2 19.1 54.8 20.4 57.6 35.1
1.7 19.8 23.9 0.302 0.0 25.8 49.4 54.0 32.8 18.4 54.2 20.7 57.6 34.8
0.5 7.5 25.2 0.396 0.0 25.7 50.5 51.0 29.1 19.1 52.5 19.1 40.0 31.9
0.8 11.4 30.2 0.992 0.0 26.6 49.8 55.0 34.7 21.1 55.0 21.9 32.4 32.9
+ ours 2.7B 4.2 47.7 49.4 3.793 58.6 41.2 55.5 76.0 59.3 28.8 72.6 21.7 57.6 52.4
5.0 48.3 53.2 3.577 60.4 41.0 55.6 70.0 59.6 25.8 72.7 21.7 57.2 51.6
4.7 43.2 50.6 3.416 60.2 41.0 55.6 74.0 60.4 27.8 72.8 21.7 57.2 52.3
4.4 48.4 55.4 3.920 60.2 40.8 56.1 73.0 59.1 27.8 72.9 21.5 57.2 52.1
4.4 53.9 54.7 3.922 61.4 40.8 56.7 73.0 60.2 27.4 73.0 21.5 57.4 52.4
Table 4: Full Results of 5 individual runs (continued).
Model # WoW ED BST WoI Avg. Pile Wikitext GPT Epoch
Params (F1) (F1) (F1) (F1) (F1) (PPL) (PPL) $\uparrow$
Neo 125M 10.5 8.4 9.7 11.3 10.0 20.1 38.0 3.75 -
+ KUMPR 125M 5.8 6.0 6.1 6.9 6.2 444.7 2396.7 1.01 3
1.1 1.1 0.7 1.0 1.0 200.8 437.8 1.20 2
0.7 1.5 0.5 0.1 0.7 >10000 >10000 0.93 5
4.7 6.4 5.9 5.5 5.6 2778.3 >10000 0.99 3
4.3 4.7 4.7 5.8 4.9 1212.3 9470.2 1.12 3
+ ours 125M 11.6 9.0 9.9 11.7 10.6 21.5 39.2 3.48 19
10.9 8.4 9.6 11.1 10.0 22.0 40.7 3.68 22
11.0 8.4 9.0 11.2 9.9 21.6 40.4 3.98 24
11.3 9.0 10.0 11.2 10.4 21.4 39.7 4.27 21
11.3 8.7 9.6 11.7 10.4 21.6 40.4 4.20 21
Neo 1.3B 12.7 10.4 12.1 13.8 12.3 13.2 18.7 2.72 -
+ KUMPR 1.3B 0.0 0.0 0.0 0.0 0.0 >10000 >10000 2.68 1
0.0 0.1 0.0 0.0 0.0 >10000 >10000 0.95 3
0.0 0.0 0.0 0.1 0.0 >10000 >10000 0.88 2
0.1 0.0 0.8 0.2 0.1 >10000 >10000 0.86 2
0.3 0.0 0.8 1.1 0.6 2164.8 >10000 0.93 2
+ ours 1.3B 12.4 10.3 11.7 13.0 11.9 14.1 19.3 4.35 28
12.1 10.0 11.7 13.5 11.8 14.1 19.3 4.72 29
12.8 10.5 12.1 13.4 12.2 14.3 19.4 4.66 30
12.6 10.4 12.0 13.9 12.2 14.2 19.3 2.96 33
12.6 11.1 12.2 13.9 12.5 14.1 19.2 4.96 26
Neo 2.7B 12.5 10.8 12.4 13.6 12.3 12.0 16.2 3.65 -
+ KUMPR 2.7B 10.9 3.2 9.7 11.6 8.8 39.2 79.9 1.02 7
10.6 0.8 9.7 11.9 8.2 66.4 166.8 1.00 8
4.9 0.1 9.0 9.3 5.8 111.2 370.8 0.94 6
0.1 0.0 0.3 0.4 0.2 >10000 >10000 0.96 6
3.5 3.2 3.8 4.3 3.7 227.1 398.6 1.08 6
+ ours 2.7B 12.1 10.4 12.2 13.0 11.9 13.0 16.9 3.70 29
12.3 10.3 12.3 12.8 12.0 13.2 16.9 4.38 34
12.2 10.0 11.7 13.1 11.7 13.3 17.0 4.13 38
12.8 10.9 12.9 13.5 12.5 13.1 17.1 5.29 34
12.3 11.2 12.2 13.3 12.2 13.1 17.0 4.51 28
Appendix B Learning Rate

### v2

@@ 3.6 Training
${}\mathcal{L}=\mathcal{L}_{fgt}+\alpha\mathcal{L}_{lrn}+\beta\mathcal{L}_{kl},$ (6)
where $\alpha,\beta>0$ are positive hyper-parameters that control the contributions of the different optimizing objectives.
@@ 4 Experiments
In this section, we first describe the datasets used for training and evaluation, as well as the metrics employed to assess performance. Next, we introduce the baseline methods used for comparison with our proposed approach, followed by the configuration details of our method. Finally, we present and analyze the experimental results.
@@ 4.1 Datasets
To evaluate ICU’s learning and unlearning capabilities, we selected two types of datasets: Target Datasets, which assess the unlearning performance, and Downstream Dataset, which evaluates the original capabilities of the models.
Target Dataset.
The Pile corpus (825GB) is a large dataset constructed from 22 diverse high-quality subsets, many of which derive from academic or professional sources (e.g. books, open source code) (Gao et al., 2020). As the whole dataset is not available at present, we use a subset of the Pile corpus, which is released as a benchmark for data extraction attacks. [^2] https://github.com/google-research/lm-extraction-benchmark Designed to be easy-to-extract, the subset contains 15,000 samples, randomly sampled from the Pile training dataset. Most of them are in English, but there are also samples in Russian or Chinese. Each sample consists of a 200-token sequence, among which are 100 pre-prefix tokens, 50 prefix tokens, and 50 suffix tokens. Following Jang et al. (2023), we only use the prefix and suffix tokens in our experiments.
Downstream Dataset.
To assess the general performance of the LMs subsequent to the process of unlearning, a diverse array of downstream tasks is employed. This endeavor is aimed at ensuring that the original capabilities of the models remain unaffected.
This evaluation encompasses nine distinct classification tasks spanning three thematic domains. Specifically, these domains include linguistic reasoning tasks such as Hellaswag (Zellers et al., 2019) and Lambada (Paperno et al., 2016), as well as assessments of commonsense reasoning through Winogrande (Sakaguchi et al., 2021) and COPA (Gordon et al., 2012). Additionally, scientific reasoning abilities are evaluated through tasks such as ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), Piqa (Bisk et al., 2020), MathQA (Amini et al., 2019), and PubmedQA (Jin et al., 2019).
Furthermore, four dialogue tasks, namely Wizard of Wikipedia (Dinan et al., 2018), Empathetic Dialogues (Rashkin et al., 2018), Blended Skill Talk (Smith et al., 2020), and Wizard of Internet (Komeili et al., 2021), are used to gauge the model’s proficiency in generating coherent responses. In addition, we measure the perplexity of the unlearned models on the validation set of Pile and Wikitext. Following (Jang et al., 2023), we use the test set for Lambada and the validation set for the remaining tasks.
@@ 4.2 Metrics
As stated in Section 4.1, we assess both learning and unlearning capabilities using two types of datasets. For evaluating unlearning performance, we follow Jang et al. (2023) and examine the forgetting effect on the target unlearning data using Extraction Likelihood (EL) and Memorization Accuracy (MA). In the work of Jang et al. (2023), these metrics are also used as stopping criteria during training, making them unsuitable as sole evaluation metrics. As mentioned in Section 3.5, we additionally employ BERTScore and BLEU to measure forgetting during the Iterative Unlearning Refinement (IUR) module.
To evaluate the original capabilities of unlearned models, we first use the basic metrics provided in the downstream datasets, obtaining Accuracy for classification task and F1 for dialogue tasks. Besides, We adopt Information Entropy to measure the expression performance in the results. Furthermore, we also employ GPT-4 and Human Evaluation to assess the text generated by the unlearned models. The following details the metrics used:
Extraction Likelihood (EL) is introduced by Jang et al. (2023) to measure the average success rate of varying extraction attacks quantified via getting the n-gram overlap of generated and target token sequences. It is computed by the following equation:
$\mathrm{EL}_{n}(\bm{x})=\frac{\sum_{t=1}^{T-n}\mathrm{OVERLAP}_{n}(f_{\theta}(x_{<t}),x_{\geq t})}{T-n},$ (7)
$\mathrm{OVERLAP}_{n}(\bm{a},\bm{b})=\frac{\sum_{c\in ng(\bm{a})}\mathbbm{1}\{c\in ng(\bm{b})\}}{|ng(\bm{a})|}.$ (8)
Memorization Accuracy (MA) (Tirumala et al., 2022) quantifies how much model $f_{\theta}$ has memorized the given token sequences, which is defined as follows:
$\mathrm{MA}(\bm{x})=\frac{\sum_{t=1}^{T-1}\mathbbm{1}\{\mathrm{argmax}(P_{\theta}(\cdot|x_{<t})=x_{t}\}}{T-1}.$ (9)
@@
$H(x)=-\sum_{i=1}^{n}p_{i}\log_{2}p_{i}.$ (10)
GPT Evaluation uses GPT-4 (Achiam et al., 2023) to evaluate the unlearned models in two perspectives: whether the model generates text without prior knowledge of key information in the referenced target data and whether the generated sequences are coherent. The prompts can be found in Appendix A.
Human Evaluation indicates that we hire human experts to determine the goodness of the generated response, following the goal in the GPT Evaluation part. The annotation agreement can be found in Appendix B.
@@ 4.3 Baseline methods
Our experiments use the GPT-Neo model family (125M, 1.3B, 2.7B) (Black et al., 2021), which is pre-trained on the Pile corpus. Following Jang et al. (2023), we utilize the Opt model family (125M, 1.3B, 2.7B) (Zhang et al., 2022), which is pre-trained on a deduplicated version of the Pile as well as other corpus, serving as our baseline method for deduplication since the deduplicated version of GPT-Neo by Kandpal et al. (2022) is not publicly available. For the approximate unlearning methods, we include KUMPR (Jang et al., 2023), DPO (Maini et al., 2024), KL (Maini et al., 2024) and LLMU (Yao et al., 2023) as other baseline methods on GPT-Neo models to show the effectiveness of our proposed method. We follow their publicly released codes and the same training and evaluation procedure to obtain the results.
@@ 4.4 Configurations
For each model size (125M, 1.3B, 2.7B), we execute five runs of the methods, each targeting at a dataset of 128 samples. In the Contrastive Learning Enhancement module (Section 3.4), we utilize the remaining Pile subset for the Analogous Data Construction, and all-MiniLM-L6-v2 model to conduct KNN sampling. The model is optimized by Adam (Kingma & Ba, 2014) with a learning rate of $65e-$ 6, and $\alpha=0.5,\beta=1.0$. We regard the model to have “forgotten” the target dataset with an average of $EL_{10}(\bm{x})<0.0499$ and $MA(\bm{x})<0.5994$. The filtering thresholds during iteration are $a=0.3$ and $b=0.01$ as introduced in Section 3.5.
@@
Table 1: Main Results showing the average of 5 random samples. Cls Avg. denotes the average accuracy of the 9 classification datasets, and Dia Avg. denotes the average F1 score of the 4 dialogue datasets. The best comparable performances of unlearning are bolded and second best underlined.
@@ 4.5 Main results
The results of all methods are summarized in Table 1. Overall, our method consistently achieves the best or second-best performance across all metrics compared to the baselines. Although methods like KUMPR, DPO, KL and LLMU may exhibit seemingly better unlearning ability measured by EL, MA, and BERT, our method preserves the generation ability of the model. Compared to Opt, our model unlearns the original model better. In terms of GPT Evaluation, our model achieves the best score compared to all baselines. Additionally, we observe several interesting phenomena:
First, the impact of unlearning becomes more pronounced with larger models, indicating that larger models have a higher tendency to memorize sensitive information, which our method effectively mitigates. Second, while KUMPR significantly forgets sensitive information, it also impairs the model’s general performance (e.g., PPL). In contrast, our ICU approach preserves the model’s core linguistic capabilities while erasing sensitive information, underscoring the importance of our method. Full results for each sample are provided in Appendix G.
@@ 4.6 Ablation study
As discussed in Section 3.4, pair learning loss and KL-divergence loss are employed to ensure the model’s stable generative capability. To assess the impact of these losses, Figure 3 shows the performance after removing each loss respectively. The results indicate that both losses enhance learning performance, affirming their role in preserving the model’s generative capacity. Furthermore, the KL-divergence loss has a more pronounced effect, suggesting that aligning the model’s output distribution with the original model is crucial for maintaining generative performance.
Figure 3: Ablation results on GPT-Neo 125M.
4.7 Parameter sensitivity
@@
4.8 Case study
Figure 4: Case study for the comparison between various methods.
@@ 4.8 Case study
To provide a clearer comparison of our methods, we present a case study demonstrating the balance between learning and unlearning. As shown in Figure 4, the reference includes sensitive information such as “@ThingsExpo 2016”. Before unlearning, the original models (e.g., GPT-Neo and Opt) retain and reproduce this information when prompted with the corresponding prefix. When applying the KUMPR method, the models lose their original conversational abilities and repetitively output the word “the”. DPO method also produces sentences not coherent. KL and LLMU methods both repeat same sentence. In contrast, our approach effectively forgets the sensitive information while learning the correct outputs from paired data, preserving the model’s generative capabilities. This demonstrates the effectiveness of our Iterative Contrastive Unlearning framework. Additional cases are provided in Appendix F.
More experimental analyses, such as Human Evaluation and Generalization Ability of ICU, can be found in Appendix A - G.
5 Conclusion and future work
@@
In this paper, we explored the direction of machine unlearning for Generative Language Models. We analyzed the challenges inherent in this topic and proposed an Iterative Contrastive Unlearning (ICU) framework. Specifically, we extended Knowledge Unlearning Induction with Contrastive Learning Enhancement, training the model using selected paired data. Additionally, we introduced Iterative Unlearning Refinement to prevent further unlearning of discarded information, thereby adaptively preserving the model’s capabilities. Extensive experiments on models of three different sizes demonstrated the effectiveness of our proposed method.
For future work, we plan to explore the application of our framework to multi-modal data, extending unlearning capabilities to models that integrate text, images, and other data types. Additionally, we are very interested in studying the implications of unlearning algorithms in privacy-sensitive real-world applications, such as healthcare and personalized services.
Appendix A GPT prompts
The prompts we employed in the evaluation part (Section 4.2) is as follows:
Final score: 10
Final score: 0
Final score: 2
Now it’s your turn:
Appendix B Human evaluation
To enhance the credibility of the GPT-4 grading, we also include human annotation results. Specifically, for each sample, human evaluators rated the text generated by the models on a scale from 1 (low quality or very similar to reference) to 8 (high quality and different from reference). The annotators are only presented with only the prefix, referenced suffix and the generated text by the model and are unaware of the full scope of the work.
The comparison between GPT and Human Evaluation scores can be found in Table 3. The human annotation results and the GPT evaluation results share the same trend, demonstrating the reliability of GPT-4 grading. Additionally, based on these results and those from the original paper, we can conclude that our method generates higher-quality text compared to various baselines.
Table 3: Human annotation results on different methods.(Avg. Human scores / Avg. GPT scores.)
# params Neo Opt Neo + KUMPR Neo + ICU (ours)
125M 3.1 / 3.75 3.7 / 3.08 1.1 / 1.05 5.7 / 3.92
1.3B 1.8 / 2.72 3.1 / 3.63 1.4 / 1.26 4.8 / 4.33
2.7B 1.5 / 2.11 3.2 / 3.65 1.1 / 1.00 4.7 / 4.40
Appendix C Generalization ability of ICU
In this section, we discuss the generalization ability of our proposed Iterative Contrastive Unlearning framework to LLMs other than the GPT-Neo model. As introduced in Section 3, our proposed ICU framework for generative LMs, which can be applied to various advanced LLMs, such as Llama, Bloom, etc. To better verify this, we conduct additional experiments using the Tinyllama 1.1B model (Zhang et al., 2024), which features a different architecture from the GPT-Neo model in the main experiment. Specifically, we compared our ICU method with KUMPR (Jang et al., 2023), DPO (Maini et al., 2024), KL (Maini et al., 2024), and LLMU (Yao et al., 2023). The results are summarized in Table 4. Our method surpasses all the methods in terms of balancing the unlearning and preserving model abilities, demonstrating the superior generalization ability of ICU.
Table 4: Comparisons of different methods on TinyLlama 1.1B.
Model EL10 MA BERT Entropy Cls Avg. Dia Avg. Pile Wikitext GPT Epoch
(%) $\downarrow$ (%) $\downarrow$ (F1) $\downarrow$ $\uparrow$ (ACC) $\uparrow$ (F1) $\uparrow$ (PPL) $\downarrow$ (PPL) $\downarrow$ $\uparrow$
TinyLlama 56.2 76.8 71.2 4.742 46.2 12.4 12.8 10.7 - -
TinyLlama + KUMPR 0.0 0.3 21.4 2.015 34.9 0.0 >10000 2152.2 1.99 4.0
TinyLlama + DPO 3.1 52.6 38.3 1.392 40.8 11.7 13.9 11.5 1.82 3.0
TinyLlama + KL 1.8 52.2 31.9 2.753 42.8 12.3 36.7 14.0 1.63 6.0
TinyLlama + LLMU 2.5 46.0 41.9 3.683 45.6 12.4 14.9 11.2 2.82 8.0
TinyLlama + ICU (ours) 1.4 44.0 44.8 3.932 45.8 12.3 14.3 11.0 3.35 10.0
Appendix D Learning rate analysis